OpenStudy (anonymous):

ASAP Please Eliminate the "z" variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1 4x - 3y = -14 4x + y = -14 6x + 5y = 13 6x - 5y = -13

OpenStudy (anonymous):

@yanee U KNOW TISS

OpenStudy (anonymous):

multiply second by negative one x+3y+2z=-16 -2x+y-2z=15 just combine -x+4y=-1 So in quick terms, 2x - 2y - z = 1 would be your answer, hope this helped